the speed of the boat  in still water is 24 km/h and speed of the stream is 41cm/h.the time taken by the boat to travel from a to b downstream is 36 minute less than time taken by the same boat to travel from b to c upstream. if the distance between a and b is 4 km more than the distance betwee b and c ,what is the disance between a and b?

Grade:12th pass

1 Answers

Eshan
5 years ago
Dear student,

Speed of the boat upstream=41km/h-24km/h=17km/h

Speed of the boat downstream=41km/h+24km/h=65km/h
Let the distance between a and b be ‘s’ and time taken to cover it be ‘t’.

Hence s=65t

It is given that\dfrac{s-4}{t+\dfrac{36}{60}}=24
\implies s=29.17km